3) There are 1,000 identical perfectly competitive real-estate firms selling office space in Syracuse, NY. The Marginal Cost of producing each square foot of space is constant and equal to $20. There are no fixed costs of production. So the firm’s short-run and long-run cost function is c(q) = 20q. The market demand is Q = 10,000 - 250p.

a) What is the equilibrium price and quantity of office space in the real-estate market in the short-run? How much does each firm sell at this equilibrium?

b) What is the profit of each firm? What is the total profit of all 1,000 firms?

c) What is the consumer surplus

d) Calculate the value of total surplus at the market equilibrium?
Answer #1

TC=20g mc- altc) - 20 da Maket At equilibrium, P=me P=20 askeat quantity Q = 10000 - 250p Q = 10000 – 250 (20) -5ooo Each 1oo